OpenStudy (anonymous):

Find the first five terms given the explicit formula :an=-11+7n

OpenStudy (mathstudent55):

Plug in 1 through 5 in for n, and evaluate the expression.

OpenStudy (anonymous):

18, 25 add 7 to 11, 5 times.

OpenStudy (mathstudent55):

\(a_1 = -11 + 7(1) \) \(a_1 = -11 + 7\) \(a_1 = -4\)